# Open Packing in Graphs: Bounds and Complexity

### v1

@@ 3.1 $K_{1,3}$-free Graphs
Note that $V(G^{\prime})=V(G)\cup E(G)\cup A$. We construct $S^{*}$ from $S$ in two phases. First, we construct an open packing $S_{1}$ in $G^{\prime}$ such that $S_{1}\subseteq V(G)\cup E(G)$ and $|S_{1}|=|S|$. If $S\cap A=\emptyset$, then $S_{1}:=S$. Else $S\cap A\neq\emptyset$ and let $S\cap A=\{v_{i_{1}},v_{i_{2}},\ldots,v_{i_{p}}\}$. Then, $S_{1}:=(S\setminus\{v_{i_{1}},v_{i_{2}},\ldots,v_{i_{p}}\})\cup\{u_{i_{1}},u_{i_{2}},\ldots,u_{i_{p}}\}$. By repeated application of Claim 2, $S_{1}$ is an open packing in $G^{\prime}$ such that $|S_{1}|=|S|$ and $S_{1}\subseteq V(G)\cup E(G)$. Next, we construct an open packing $S^{*}$ of $G^{\prime}$ from $S_{1}$ such that $|S^{*}|=|S_{1}|$ and $S^{*}\subseteq V(G)$. If $S_{1}\cap E(G)=\emptyset$, then $S^{*}:=S_{1}$. If not, let $S_{1}\cap E(G)=\{e_{1},e_{2},\ldots,e_{q}\}$ and let $u_{\ell_{i}}$ be an end vertex of $e_{i}$ in $G$ for $1\leq i\leq q$. Note that for $1\leq i<j\leq q$, $u_{\ell_{i}}\neq u_{\ell_{j}}$, otherwise $\{u_{\ell_{i}},e_{i},e_{j}\}$ induces a triangle (by Equation (1)) in $G^{\prime}$ with $e_{i},e_{j}\in S_{1}$ (an open packing in $G^{\prime}$), a contradiction to Observation 2. Thus, $S^{*}:=(S_{1}\setminus\{e_{1},e_{2},\ldots,e_{q}\})\cup\{u_{\ell_{1}},u_{\ell_{2}},\ldots,u_{\ell_{q}}\}$. By repeated application of Claim 4, $S^{*}$ is an open packing in $G^{\prime}$ with $|S^{*}|=|S_{1}|=|S|=k$ and $S^{*}\subseteq V(G)$. By Claim 1, $S^{*}$ is an independent set in $G$. This proves Claim 5.
@@
Max-Independent Set cannot be approximated within a factor of $n^{(1-\epsilon)}$ for any $\epsilon>0$, in general graphs unless NP=P.
@@
Max-Open Packing is hard to approximate within a factor of $N^{\frac{1}{2}-\delta}$ for any $\delta>0$ in $K_{1,3}$-free graphs unless $NP=P$, where $N$ denotes the number of vertices in a $K_{1,3}$-free graph.
@@
$|S|$ $>\cdot\frac{\alpha(G)}{n^{1-2\delta}}$
$|U|$ $>\cdot\frac{\alpha(G)}{n^{1-2\delta}}$
@@
The above inequality implies that Max-Independent Set has a $(n^{1-\epsilon})$-factor approximation algorithm for some $\epsilon>0$ which contradicts Theorem 4. ∎
Next, we give a parameterized intractability result for the problem Open Packing on $K_{1,3}$-free split graphs using Construction 1 and the following result from the literature.
@@
Note that Independent Set parameterized by solution size is W[1]-complete [4]. Lemma 1 implies that every instance $(G,k)$ of Open Packing on $K_{1,3}$-free graphs can be (FPT) reduced into an instance $(G^{[o]},k)$ of Independent Set, and hence Open Packing parameterized by solution size on $K_{1,3}$-free graphs is in W[1]. The guarantee of Construction 1 implies that every instance $(G,k)$ of Independent Set can be (FPT) reduced into an instance $(G^{\prime},k)$ of Open Packing on $K_{1,3}$-free graphs, and thus Open Packing parameterized by solution size is W[1]-hard on $K_{1,3}$-free graphs. ∎
@@ 3.2 Subclasses of $(P_{4}\cup rK_{1})$-free Graphs
In this section, we prove bounds on the total domination number and the open packing number in the class of (i) $(P_{4}\cup rK_{1})$-free graphs, (ii) $(P_{3}\cup rK_{1})$-free graphs, (iii) $(K_{2}\cup rK_{1})$-free graphs and (iv) $rK_{1}$-free graphs. These bounds eventually lead to solve Total Dominating Set and Open Packing in polynomial time in the aforementioned graph classes. The following lemmas are the tools used to connect the bounds with polynomial time algorithms.
@@
Given a graph $G$ and a vertex subset $D$ of $G$, it can be tested whether $D$ is a total dominating set of $G$ or not in $O(n+m)$-time.
@@
Let $G\in\mathcal{G}$ and let $\mathcal{D}_{\ell}$ be the set of all total dominating sets of size at most $\ell$ in $G$. Also, let $D$ be an optimal total dominating set in $G$. Then, since $|D|=\gamma_{t}(G)\leq\ell$, $D\in\mathcal{D}_{\ell}$. Also, $\mathcal{D}_{\ell}\subseteq\mathcal{V}_{\ell}=\{X\subseteq V(G)\,:\,|X|\leq\ell\}$. Note that the collection $\mathcal{V}_{\ell}$ can be found in $O(n^{\ell})$ time. Also, for every $X\in\mathcal{V}_{\ell}$, it can be tested whether $X\in\mathcal{D}_{\ell}$ in $O(n+m)$ time using Lemma 2. This together with the fact that $|\mathcal{V}_{\ell}|=O(n^{\ell})$ imply that the collection $\mathcal{D}_{\ell}$ can be found in $O(n^{\ell}(n+m))$ time. Since every $D^{\prime}\in\mathcal{D}_{\ell}$ is a total dominating set $\gamma_{t}(G)\leq\min\{|D^{\prime}|\,:\,D^{\prime}\in\mathcal{D}_{\ell}\}$. Further, since $D\in\mathcal{D}_{\ell}$, $\gamma_{t}(G)=|D|\geq\min\{|D^{\prime}|\,:\,D^{\prime}\in\mathcal{D}_{\ell}\}$. So, $\gamma_{t}(G)=\min\{|D^{\prime}|\,:\,D^{\prime}\in\mathcal{D}_{\ell}\}$. Since $\mathcal{D}_{\ell}$ can be generated in $O(n^{\ell}(n+m))$ time, so is $\gamma_{t}(G)$ and an optimal total dominating set of $G$. ∎
The following lemma is helpful in proving a result similar to that of the above lemma for open packing.
@@
Given a graph class $\mathcal{G}$, if there exists $k\in\mathbb{N}$ such that $\rho^{o}(G)\leq k$ for every $G\in\mathcal{G}$, then (i) $G$ contains at most $O(n^{k})$ open packings and (ii) all open packings in $G$ can be computed in $O(n^{k+1})$ time for every $G\in\mathcal{G}$. So, $\rho^{o}(G)$ can be computed in $O(n^{k+1})$ time.
@@
Let $G\in\mathcal{G}$. Let $\mathcal{S}$ be the set of all open packings in $G$. We prove that the set $\mathcal{S}$ can be found in $O(n^{k+1})$ time. Let $\mathcal{V}_{k}=\{X\subseteq V(G)\,:\,|X|\leq k\}$. Then, since $\rho^{o}(G)\leq k$, $\mathcal{S}\subseteq\mathcal{V}_{k}$. Note that the set $\mathcal{V}_{k}$ is of size $O(n^{k})$ and can be found in $O(n^{k})$ time. Since $\mathcal{S}\subseteq\mathcal{V}_{k}$, $|\mathcal{S}|=O(n^{k})$. Also, for every vertex subset $X\in\mathcal{V}_{k}$, it can be tested in $O(n)$ time whether $X\in\mathcal{S}$ using Lemma 4. Thus, the set $\mathcal{S}$ can be found in $O(n^{k+1})$ time. Note that $\rho^{o}(G)=\max\{|S|\,:\,S\in\mathcal{S}\}$. Since, the set $\mathcal{S}$ can be found in $O(n^{k+1})$ time, $\rho^{o}(G)$ can be computed in $O(n^{k+1})$ time. ∎
Next, we study the bounds on total domination number and open packing number in the class of $(P_{t}\cup rK_{1})$-free graphs for $t\in\{1,2,3,4\}$ and every non-negative integer $r\geq 3-t$.
@@
In this section, we prove that for $r\geq 1$, $\gamma_{t}(G)\leq 2r+2$ and $\rho^{o}(G)\leq 2r+1$ in a connected ($P_{4}\cup rK_{1}$)-free graph $G$ (the bound on the open packing number of $(P_{4}\cup rK_{1})$-free graphs along with Lemma 5 will solve the sufficiency part of Theorem 2). Further, we show that the bound on the open packing number in these graph classes are tight (see Remark 3) and that the bound on the total domination number is tight for the class of $(P_{4}\cup K_{1})$-free graphs (see Remark 2). The following known lemma solves the problems for $P_{4}$-free graphs.
@@
The bound given in Lemma 7 is tight for the case $r=1$, and a cycle on six vertices ($C_{6}$) is a connected $(P_{4}\cup K_{1})$-free graph that satisfies this bound. Also, the graphs obtained by replacing every vertex of $C_{6}$ by a complete graph will remain as a ($P_{4}\cup K_{1}$)-free graph with total domination number four. For every $r>1$, there exists a $(P_{4}\cup rK_{1})$-free graph $G_{r}$ with $\gamma_{t}(G_{r})=2r+1$ (see Remark 3 and Fig. 4), and so the bound given in Lemma 7 is a near-optimal upper bound.
The following corollary follows from Lemmas 6 and 7 and the fact that for every $r\geq 1$, every $(K_{2}\cup rK_{1})$-free graph is a ($P_{4}\cup(r-1)K_{1}$)-free graph.
@@
The following observation that follows from the definition of open packing is helpful to prove Lemma 8.
@@
Let $S$ be an open packing in a graph $G$. Then, $G[S]$ is $\{P_{3},K_{3}\}$-free i.e., $G[S]$ is a union of $K_{2}$’s and $K_{1}$’s.
The next lemma proves the bound on the open packing number in the class of $(P_{4}\cup rK_{1})$-free graphs.
@@
We use induction on $r\geq 1$ to prove this lemma as well. Firstly, we prove the bound for the case $r=1$. Let $G$ be a connected $(P_{4}\cup K_{1})$-free graph. Then, $\rho^{o}(G)\leq 2<2r+1=3$ if $G$ is $P_{4}$-free by Lemma 6. So, assume that $G$ contains a $P_{4}$. Let $D=\{x_{1},x_{2},x_{3},x_{4}\}$ be a vertex subset of $G$ that induces a $P_{4}$ in $G$ with $x_{i}x_{i+1}\in E(G)$ for $i=1,2,3$. Then, since $G$ is $(P_{4}\cup K_{1})$-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $\{x_{1},x_{2},x_{3},x_{4}\}$. Also, every vertex in $\{x_{1},x_{2},x_{3},x_{4}\}$ is adjacent to a vertex in $\{x_{2},x_{3}\}$. So, $V(G)=\bigcup\limits_{i=1}^{4}N(x_{i})$. Thus, for any open packing $S$ in $G$, $S=S\cap V(G)=\bigcup\limits_{i=1}^{4}(S\cap N(x_{i}))$. Since $S$ is an open packing, $|S\cap N(x_{i})|\leq 1$ for $i=1,2,3,4$. Thus, $|S|\leq\sum\limits_{i=1}^{4}|S\cap N(x_{i})|\leq 4$. So, to prove the claim, it is enough to disprove the possibility of an open packing of size four in $G$. On the contrary, assume that there exists an open packing $S$ of size four in $G$. Then, by the above arguments, we can conclude that $|S\cap N(x_{i})|=1$ for every $i=1,2,3,4$. Let $S\cap N(x_{i})=\{z_{i}\}$. So, $S=\{z_{1},z_{2},z_{3},z_{4}\}$. This together with the assumption that $|S|=4$ implies $z_{i}\neq z_{j}$ (and so $x_{i}z_{j}\notin E(G)$) for $i,j\in\{1,2,3,4\}$ and $i\neq j$. Note that $z_{1}x_{1}\in E(G)$ (i.e., $z_{1}\neq x_{1}$) whereas $x_{1}x_{3},x_{1}x_{4}\notin E(G)$, so $z_{1}\notin\{x_{1},x_{3},x_{4}\}$. Further, $z_{1}\neq x_{2}$ since $x_{2}x_{3}\in E(G)$ and $z_{1}x_{3}\notin E(G)$. Similarly, it can be proved that $z_{4}\notin\{x_{1},x_{2},x_{3},x_{4}\}$. Therefore, $z_{1}z_{4}\in E(G)$, else $\{z_{1},x_{1},x_{2},x_{3}\}\cup\{z_{4}\}$ induces a $(P_{4}\cup K_{1})$ in $G$, a contradiction. Since $z_{1}z_{4}\in E(G)$, we have $z_{1}z_{2},z_{2}z_{4}\notin E(G)$ because of Observation 3. This together with the fact that $z_{1}x_{1},z_{4}x_{4}\in E(G)$ implies that $z_{2}\notin\{x_{1},x_{4}\}$. Further, since $x_{1}z_{2},z_{2}x_{4}\notin E(G)$, the vertex subset $\{x_{1},z_{1},z_{4},x_{4}\}\cup\{z_{2}\}$ induces a $(P_{4}\cup K_{1})$ in $G$, a contradiction. So $\rho^{o}(G)\leq 3$. Next, we prove the case $r\geq 2$.

### v2

@@ 3.1 $K_{1,3}$-free Graphs
Note that $V(G^{\prime})=V(G)\cup E(G)\cup A$. We construct $S^{*}$ from $S$ in two phases. First, we construct an open packing $S_{1}$ in $G^{\prime}$ such that $S_{1}\subseteq V(G)\cup E(G)$ and $|S_{1}|=|S|$. If $S\cap A=\emptyset$, then $S_{1}:=S$. Else $S\cap A\neq\emptyset$ and let $S\cap A=\{v_{i_{1}},v_{i_{2}},\ldots,v_{i_{p}}\}$. Then, $S_{1}:=(S\setminus\{v_{i_{1}},v_{i_{2}},\ldots,v_{i_{p}}\})\cup\{u_{i_{1}},u_{i_{2}},\ldots,u_{i_{p}}\}$. By repeated application of Claim 2, $S_{1}$ is an open packing in $G^{\prime}$ such that $|S_{1}|=|S|$ and $S_{1}\subseteq V(G)\cup E(G)$. Next, we construct an open packing $S^{*}$ of $G^{\prime}$ from $S_{1}$ such that $|S^{*}|=|S_{1}|$ and $S^{*}\subseteq V(G)$. If $S_{1}\cap E(G)=\emptyset$, then $S^{*}:=S_{1}$. If not, let $S_{1}\cap E(G)=\{e_{1},e_{2},\ldots,e_{q}\}$, and let $u_{\ell_{i}}$ be an end vertex of $e_{i}$ in $G$ for $1\leq i\leq q$. Note that for $1\leq i<j\leq q$, $u_{\ell_{i}}\neq u_{\ell_{j}}$; otherwise, $\{u_{\ell_{i}},e_{i},e_{j}\}$ induces a triangle (by Equation (1)) in $G^{\prime}$ with $e_{i},e_{j}\in S_{1}$ (an open packing in $G^{\prime}$), a contradiction to Observation 2. Thus, $S^{*}:=(S_{1}\setminus\{e_{1},e_{2},\ldots,e_{q}\})\cup\{u_{\ell_{1}},u_{\ell_{2}},\ldots,u_{\ell_{q}}\}$. By repeated application of Claim 4, $S^{*}$ is an open packing in $G^{\prime}$ with $|S^{*}|=|S_{1}|=|S|=k$ and $S^{*}\subseteq V(G)$. By Claim 1, $S^{*}$ is an independent set in $G$. This proves Claim 5.
@@
Max-Independent Set cannot be approximated within a factor of $n^{(1-\epsilon)}$ for any $\epsilon>0$, in general graphs unless P = NP.
@@
Max-Open Packing is hard to approximate within a factor of $N^{\frac{1}{2}-\delta}$ for any $\delta>0$ in $K_{1,3}$-free graphs unless P = NP, where $N$ denotes the number of vertices in a $K_{1,3}$-free graph.
@@
$|S|$ $>\frac{\alpha(G)}{n^{1-2\delta}}$
$|U|$ $>\frac{\alpha(G)}{n^{1-2\delta}}$
@@
The above inequality implies that Max-Independent Set has a $(n^{1-\epsilon})$-factor approximation algorithm for some $\epsilon>0$, which contradicts Theorem 4. ∎
Next, we give a parameterized intractability result for Open Packing on $K_{1,3}$-free split graphs using Construction 1 and the following result from the literature.
@@
Note that Independent Set parameterized by solution size is W[1]-complete [4]. Lemma 1 implies that every instance $(G,k)$ of Open Packing on $K_{1,3}$-free graphs can be (FPT) reduced into an instance $(G^{[o]},k)$ of Independent Set. Hence, Open Packing parameterized by solution size on $K_{1,3}$-free graphs is in W[1]. The guarantee of Construction 1 implies that every instance $(G,k)$ of Independent Set can be (FPT) reduced into an instance $(G^{\prime},k)$ of Open Packing on $K_{1,3}$-free graphs. Thus, Open Packing parameterized by solution size is W[1]-hard on $K_{1,3}$-free graphs. ∎
@@ 3.2 Subclasses of $(P_{4}\cup rK_{1})$-free Graphs
In this section, we prove bounds on the total domination number and the open packing number in the classes of (i) $(P_{4}\cup rK_{1})$-free graphs, (ii) $(P_{3}\cup rK_{1})$-free graphs, (iii) $(K_{2}\cup rK_{1})$-free graphs, and (iv) $rK_{1}$-free graphs. These bounds eventually solve Total Dominating Set and Open Packing in polynomial time in the graph classes mentioned above. The following lemmas are the tools used to connect the bounds with polynomial time algorithms.
@@
Given a graph $G$ and a vertex subset $D$ of $G$, it can be tested whether $D$ is a total dominating set of $G$ or not in $O(n+m)$ time.
@@
Let $G\in\mathcal{G}$, and let $\mathcal{D}_{\ell}$ be the set of all total dominating sets of size at most $\ell$ in $G$. Also, let $D$ be an optimal total dominating set in $G$. Then, since $|D|=\gamma_{t}(G)\leq\ell$, $D\in\mathcal{D}_{\ell}$. Also, $\mathcal{D}_{\ell}\subseteq\mathcal{V}_{\ell}=\{X\subseteq V(G)\,:\,|X|\leq\ell\}$. Note that the collection $\mathcal{V}_{\ell}$ can be found in $O(n^{\ell})$ time. Also, for every $X\in\mathcal{V}_{\ell}$, it can be tested whether $X\in\mathcal{D}_{\ell}$ in $O(n+m)$ time using Lemma 2. This, together with the fact that $|\mathcal{V}_{\ell}|=O(n^{\ell})$, implies that the collection $\mathcal{D}_{\ell}$ can be found in $O(n^{\ell}(n+m))$ time. Since every $D^{\prime}\in\mathcal{D}_{\ell}$ is a total dominating set, $\gamma_{t}(G)\leq\min\{|D^{\prime}|\,:\,D^{\prime}\in\mathcal{D}_{\ell}\}$. Further, since $D\in\mathcal{D}_{\ell}$, $\gamma_{t}(G)=|D|\geq\min\{|D^{\prime}|\,:\,D^{\prime}\in\mathcal{D}_{\ell}\}$. So, $\gamma_{t}(G)=\min\{|D^{\prime}|\,:\,D^{\prime}\in\mathcal{D}_{\ell}\}$. Since $\mathcal{D}_{\ell}$ can be generated in $O(n^{\ell}(n+m))$ time, so is $\gamma_{t}(G)$ and an optimal total dominating set of $G$. ∎
The following lemma is helpful in proving a result similar to the above lemma for open packing.
@@
Given a graph class $\mathcal{G}$, if there exists $k\in\mathbb{N}$ such that $\rho^{o}(G)\leq k$ for every $G\in\mathcal{G}$, then (i) $G$ contains at most $O(n^{k})$ open packings, and (ii) all open packings in $G$ can be computed in $O(n^{k+1})$ time for every $G\in\mathcal{G}$. So, $\rho^{o}(G)$ can be computed in $O(n^{k+1})$ time.
@@
Let $G\in\mathcal{G}$. Let $\mathcal{S}$ be the set of all open packings in $G$. We prove that the set $\mathcal{S}$ can be found in $O(n^{k+1})$ time. Let $\mathcal{V}_{k}=\{X\subseteq V(G)\,:\,|X|\leq k\}$. Then, since $\rho^{o}(G)\leq k$, $\mathcal{S}\subseteq\mathcal{V}_{k}$. Note that the set $\mathcal{V}_{k}$ is of size $O(n^{k})$ and can be found in $O(n^{k})$ time. Since $\mathcal{S}\subseteq\mathcal{V}_{k}$, $|\mathcal{S}|=O(n^{k})$. Also, for every vertex subset $X\in\mathcal{V}_{k}$, it can be tested in $O(n)$ time whether $X\in\mathcal{S}$ using Lemma 4. Thus, the set $\mathcal{S}$ can be found in $O(n^{k+1})$ time. Note that $\rho^{o}(G)=\max\{|S|\,:\,S\in\mathcal{S}\}$. Since the set $\mathcal{S}$ can be found in $O(n^{k+1})$ time, $\rho^{o}(G)$ can be computed in $O(n^{k+1})$ time. ∎
Next, we study the bounds on the total domination number and the open packing number in the class of $(P_{t}\cup rK_{1})$-free graphs for $t\in\{1,2,3,4\}$ and every non-negative integer $r\geq 3-t$.
@@
In this section, we prove that for $r\geq 1$, $\gamma_{t}(G)\leq 2r+2$, and $\rho^{o}(G)\leq 2r+1$ in a connected ($P_{4}\cup rK_{1}$)-free graph $G$ (the bound on the open packing number of $(P_{4}\cup rK_{1})$-free graphs along with Lemma 5 will solve the sufficiency part of Theorem 2). Further, we show that the bound on the open packing number in these graph classes is tight (see Remark 3) and that the bound on the total domination number is tight for the class of $(P_{4}\cup K_{1})$-free graphs (see Remark 2). The following known lemma solves the problems for $P_{4}$-free graphs.
@@
For $r=1$, the bound given in Lemma 7 is tight, and a cycle on six vertices ($C_{6}$) is a connected $(P_{4}\cup K_{1})$-free graph that satisfies this bound. Also, the graphs obtained by replacing every vertex of $C_{6}$ with a complete graph will remain as a ($P_{4}\cup K_{1}$)-free graph with total domination number four. For every $r>1$, there exists a $(P_{4}\cup rK_{1})$-free graph $G_{r}$ with $\gamma_{t}(G_{r})=2r+1$ (see Remark 3 and Fig. 4). Therefore, the bound given in Lemma 7 is a near-optimal upper bound.
The following corollary follows from Lemmas 6 and 7, and the fact that for every $r\geq 1$, every $(K_{2}\cup rK_{1})$-free graph is a ($P_{4}\cup(r-1)K_{1}$)-free graph.
@@
The following observation that follows from the definition of open packing is helpful in proving Lemma 8.
@@
Let $S$ be an open packing in a graph $G$. Then, $G[S]$ is $\{P_{3},K_{3}\}$-free, i.e., $G[S]$ is a union of $K_{2}$’s and $K_{1}$’s.
The following lemma proves the bound on the open packing number in the class of $(P_{4}\cup rK_{1})$-free graphs.
@@
We use induction on $r\geq 1$ to prove this lemma as well. Firstly, we prove the bound for the case $r=1$. Let $G$ be a connected $(P_{4}\cup K_{1})$-free graph. Then, $\rho^{o}(G)\leq 2<2r+1=3$ if $G$ is $P_{4}$-free by Lemma 6. So, assume that $G$ contains a $P_{4}$. Let $D=\{x_{1},x_{2},x_{3},x_{4}\}$ be a vertex subset of $G$ that induces a $P_{4}$ in $G$ with $x_{i}x_{i+1}\in E(G)$ for $i=1,2,3$. Then, since $G$ is $(P_{4}\cup K_{1})$-free, every vertex in $V(G)\setminus D$ is adjacent to some vertex in $\{x_{1},x_{2},x_{3},x_{4}\}$. Also, every vertex in $\{x_{1},x_{2},x_{3},x_{4}\}$ is adjacent to a vertex in $\{x_{2},x_{3}\}$. So, $V(G)=\bigcup\limits_{i=1}^{4}N(x_{i})$. Thus, for any open packing $S$ in $G$, $S=S\cap V(G)=\bigcup\limits_{i=1}^{4}(S\cap N(x_{i}))$. Since $S$ is an open packing, $|S\cap N(x_{i})|\leq 1$ for $i=1,2,3,4$. Thus, $|S|\leq\sum\limits_{i=1}^{4}|S\cap N(x_{i})|\leq 4$. So, to prove the claim, it is enough to disprove the possibility of an open packing of size four in $G$. On the contrary, assume that there exists an open packing $S$ of size four in $G$. Then, by the above arguments, we can conclude that $|S\cap N(x_{i})|=1$ for every $i=1,2,3,4$. Let $S\cap N(x_{i})=\{z_{i}\}$. So, $S=\{z_{1},z_{2},z_{3},z_{4}\}$. This, together with the assumption that $|S|=4$, implies $z_{i}\neq z_{j}$ (and so $x_{i}z_{j}\notin E(G)$) for $i,j\in\{1,2,3,4\}$ and $i\neq j$. Note that $z_{1}x_{1}\in E(G)$ (i.e., $z_{1}\neq x_{1}$) whereas $x_{1}x_{3},x_{1}x_{4}\notin E(G)$, so $z_{1}\notin\{x_{1},x_{3},x_{4}\}$. Further, $z_{1}\neq x_{2}$ since $x_{2}x_{3}\in E(G)$ and $z_{1}x_{3}\notin E(G)$. Similarly, it can be proved that $z_{4}\notin\{x_{1},x_{2},x_{3},x_{4}\}$. Therefore, $z_{1}z_{4}\in E(G)$, else $\{z_{1},x_{1},x_{2},x_{3}\}\cup\{z_{4}\}$ induces a $(P_{4}\cup K_{1})$ in $G$, a contradiction. Since $z_{1}z_{4}\in E(G)$, we have $z_{1}z_{2},z_{2}z_{4}\notin E(G)$ because of Observation 3. This, together with the fact that $z_{1}x_{1},z_{4}x_{4}\in E(G)$, implies that $z_{2}\notin\{x_{1},x_{4}\}$. Further, since $x_{1}z_{2},z_{2}x_{4}\notin E(G)$, the vertex subset $\{x_{1},z_{1},z_{4},x_{4}\}\cup\{z_{2}\}$ induces a $(P_{4}\cup K_{1})$ in $G$, a contradiction. So $\rho^{o}(G)\leq 3$. Next, we prove the case $r\geq 2$.